OpenStudy (anonymous):

Which equation demonstrates the opposite of a sum property? A. (22 + 10) + 3 = (10 + 22) + 3 B. –5 + 13 = 13 + (–5) C. (12 + (–4)) + 14 = 12 + ((–4) + 14) D. –(–6 + 17) = 6 + (–17)

OpenStudy (anonymous):

@calculusxy

OpenStudy (anonymous):

@1351250faith

OpenStudy (freckles):

a+b is a sum -(a+b) is read to be the opposite of the sum of a and b so the opposite of (a+b) is -(a+b)=-a-b

OpenStudy (anonymous):

sooooooo? a

OpenStudy (freckles):

I don't see anywhere in choice a where you have something like -(a+b)=-a-b

OpenStudy (anonymous):

D

OpenStudy (freckles):

try to use what I have said

OpenStudy (freckles):

yes that seems right to me

OpenStudy (anonymous):

What value of h makes the statement true? (17 + h) + (–9) = 17 + (5 + (–9)) A. –21 B. –5 C. 5 D. 21

OpenStudy (anonymous):

21